OpenStudy (anonymous):

x + 2 = y + a y - z = a - 5 z = x - 2 + a According to the system of equations above, what is the value of a?

OpenStudy (solomonzelman):

Take the first equation, and subtract 2 from both sides. x + 2 = y + a x + 2 -2 = y + a -2 x = y + a -2 substitute y + a -2 for x, into 3rd equation. z = x - 2 + a ----> z = (y + a -2) - 2 + a ---> z=y+a So now we have two equations. y - z = a - 5 z=y+a subtract y from both sides in a second equation. z-y=y-y+a z-y=a so now we have y - z = a - 5 and z-y=a multiply the second equation by -1 y-z=-a we know that y-z= -a and y-z = a - 5 so we can say that -a=a-5 subtract a from both sides, -2a=5 divide both sides by 2 a= -5/2

OpenStudy (solomonzelman):

Do you need more help understanding this?

OpenStudy (shamil98):

-a=a-5 -2a = -5 a = 5/2 not -5/2 @SolomonZelman